Required information Problem 11-4A Warranty expense and liability estimation LO P4 The following information applies to the questions displayed below. On October 29, 2016, Lobo Co. began operations by purchasing razors for resale. Lobo uses the perpetual inventory method. The razors have a 90-day warranty that requires the company to replace any nonworking razor. When a razor is returned, the company discards it and mails a new one from Merchandise Inventory to the customer. The companys cost per new razor is $20 and its retail selling price is $75 in both 2016 and 2017. The manufacturer has advised the company to expect warranty costs to equal 8% of dollar sales. The following transactions and events occurred. 2016 Nov. Sold 105 razors for $7, 875 cash. 30 Recognized warranty expense related to November sales with an adjusting entry. Dec. 9 Replaced 15 razors that were returned under the warranty 16 Sold 220 razors for $16, 500 cash. 29 Replaced 30 razors that were returned under the warranty, 31 Recognized warranty expense related to December sales with an adjusting entry 2017 Jan. 5 Sold 150 razors for $11,250 cash 17 Replaced 50 razors that were returned under the warranty. 31 Recognized warranty expense related to January sales with an adjusting entry.Problem 11-4A Part 1 1.1 Prepare journal entries to record above transactions and adjustments for 2016. View transaction list X: 1 Record the sales revenue of 105 razors for $7,875 cash. 2Record the cost of goods sold for 105 razors. Record the estimated warranty expense at 8% of November sales. 4 Record the replacement of 15 razors that were returned under the warranty. 5 Record the sales revenue of 220 razors for $16,500 cash. 6 Record the cost of goods sold for 220 razors. 7Record the replacement of 30 razors that were returned Note:-journal entry has been entered Credit Record entry Clear entry View general iournal1.2 Prepare journal entries to record above transactions and adjustments for 2017. View transaction list Record the sales revenue of 150 razors for $11,250 cash. Record the cost of goods sold for 150 razors. Record the replacement of 50 razors that were returned 1 2 3 under the warranty. 4 Record the adjusting entry for warranty expense for the month of January 2017. Credit Note journal entry has been entered Record entry Clear entry View general journal
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 7,875
Sales 7,875
11-Nov Cost of goods sold 2100 =105*20
Merchandise inventory 2,100
30-Nov Warranty expense 630 =7875*8%
Estimated warranty liability 630
9-Dec Estimated warranty liability 300 =15*20
Merchandise inventory 300
16-Dec Cash 16,500
Sales 16,500
16-Dec Cost of goods sold 4,400 =220*20
Merchandise inventory 4,400
29-Dec Estimated warranty liability 600 =30*20
Merchandise inventory 600
31-Dec Warranty expense 1,320 =16500*8%
Estimated warranty liability 1,320
1.2
Date General Journal Debit Credit
5-Jan Cash 11,250
Sales 11,250
5-Jan Cost of goods sold 3,000 =150*20
Merchandise inventory 3,000
17-Jan Estimated warranty liability 1000 =50*20
Merchandise inventory 1,000
31-Jan Warranty expense 900 =11250*8%
Estimated warranty liability 900